OpenStudy (anonymous):

a = 3(b+4) solve for b. What's the answer?

OpenStudy (geometry_hater):

distribute the 3 So we have a= 3b+ 12 now try to get b by itself

OpenStudy (anonymous):

first divide 3 on both sides and then subtract 4 on both sides to get b = (a/3) - 4

OpenStudy (anonymous):

a=3b+12 a-12=3b b=(a-12)/3 b=a/3-4